# Long Chain-of-Thought for Reasoning Large Language Models

### v1

@@ 5.2.1 Prompt-based Refinement Generation
Research on prompt-based refine generation focuses on enhancing the performance of LLMs through iterative self-refinement mechanisms [408, 762, 68, 333, 723, 539, 582]. A prominent approach involves prompting RLLMs to generate initial outputs, followed by self-feedback that iteratively refines and improves performance across tasks such as dialogue generation and mathematical reasoning [458, 375, 768, 471, 384, 240, 535], which even much reduce the hallucinations [196, 215]. Noteworthy methods, like Self-Backtracking [661], Refiner [417], and BackMath [740], allow LLMs to adjust their reasoning autonomously, reducing unnecessary complexity in decision-making [612]. Further, Havrilla et al. [164] extend the paradigm by integrating overall-level and step-level refinements, improving refinement performance. Yang et al. [664] propose a method to decompose the self-correction capability of LLMs into "confidence" and "critique" capacities, designing probabilistic metrics to evaluate them and exploring the role of reflection mechanisms in model behavior. Additionally, MCTSr [726], LLM2 Yang et al. [652], ReST-MCTS* [725] and ReARTeR [493] emphasize dynamic reflection through iterative error correction and confidence adjustments, allowing models to autonomously refine reasoning strategies [122]. He et al. [166] extend this paradigm to multi-agent scenarios, improving agent system performance. However, without oracle feedback, RLLM’s self-refinement process fails, causing instability in both intermediate and final answers, leading to biases in simple factual queries and introducing cognitive biases in complex tasks [738].
@@ 6.1.1 Vertical Scaling
Vertical scaling refers to extending the reasoning output within a single model generation, significantly boosting model performance [272]. Early works by Fu et al. [124] and Jaech et al. [208] show that increasing the length of the reasoning path can greatly improve performance. Building on this, later studies [214, 277] further explore enhancing logical depth through tree-based searches within a fixed compute budget, resulting in notable performance gains. Building upon this, Muennighoff et al. [391] introduce a test-time scaling method that improves reasoning by fine-tuning and budget forcing, yielding substantial gains with additional computing at test time. To address the constraints of attention spans, some studies focus on expanding reasoning length in latent spaces. Geiping et al. [136] and Chen et al. [77] enhance test-time reasoning performance by implicitly scaling computation in latent space through recurrent depth.
@@ 6.1.2 Parallel Scaling
Parallel scaling refers to the process of increasing the number of reasoning iterations during model generation and then verfiy these results to get the final output, which significantly enhances model performance [2, 610, 40]. Initially, Wang et al. [580] introduce the concept of self-consistency, demonstrating that multiple sampling processes followed by majority voting for effective exploration.
@@
The primary focus of recent research is optimizing verification, which can be categorized into two types: (1) Overall Verification: Recent works [783, 591] divide the scaling process into two stages: "reasoning" and "self-verification." By replacing majority voting in self-consistency with self-verification, these approaches show significant improvements [758, 59, 800]. In code scenarios, WoT [750], CISC [502] and S* [278] scale the Long CoT in parallel, using output confidence or code execution results for verification, effectively assessing reasoning quality [449, 135]. Further, Nye et al. [399] and Weir et al. [597] train RLLMs to simulate code execution, removing the need for test cases in code-related parallel scaling. Chain-of-Verification [66] introduces meta-verification, sampling multiple verification instances to identify the correct one. Kim et al. [245], Chen et al. [78], and Vacareanu et al. [535] validate this approach empirically by evaluating answer correctness based on reasoning path properties. Moreover, Li et al. [301] tune a specific RLLM to verify and aggregate answers, showing improved performance. This suggests that PRM cannot replace a specially trained RLLM for verification due to training goal biases [755]. Finally, Kang et al. [236] leverage self-uncertainty to select the best results. (2) Step Verification: Building on this, numerous researchers have explored step-level or finer-grained verification [61, 327]. Notably, DIVERSE [304], SSC-CoT [766], and Fine-grained Self-Consistency [66] combine diverse reasoning paths with step-level verification. In addition, Snell et al. [477], Wu et al. [610], Luo et al. [358], Wang et al. [552], Wu et al. [604], and Liu et al. [343] investigate how optimal scaling strategies based on MCTS can enhance smaller language models’ performance. Their findings show that a 1B RLLM can outperform a 405B model on complex tasks through parallel scaling [690]. Despite these advancements in verification, Chen et al. [66] demonstrate that these strategies cannot surpass Best-of-N methods, suggesting that breakthroughs cannot solely rely on optimization-based verification [75].
@@
Another key area of research focuses on generating diverse paths or strategies for efficient scaling [615, 548]. For instance, Zeng et al. [715] aggregate the shortest yet most varied reasoning paths for better scalability. Similarly, Du et al. [110] adjust the sampling temperature to increase diversity, leading to improved scaling. Zhang et al. [734] and Liu et al. [334] optimize both candidate solution generation (e.g., prompts, temperature, and top-p) and reward mechanisms (such as self-evaluation and reward types), offering diverse strategies for parallel scaling. Moreover, Qin et al. [435], Luo et al. [361], and Yu et al. [691] enhance RLLM reasoning by scaling sampling across multiple natural and programming languages or varied expressions. Finally, Yang et al. [660] introduces a method where a small set of seed data, with varied response lengths, guides the model to engage in deeper reasoning by selecting the shortest correct responses across various inference efforts.
@@ 6.2.1 RL Strategies
Recent advancements in RL strategies for exploration have led to notable improvements in various tasks, particularly in reasoning tasks [490, 261, 213, 378, 621].
@@
(1) Reward-model-free RL: The first series of work focuses on RL optimization algorithms. Additionally, OREO [555] propose an offline RL method that optimizes the soft Bellman equation, improving credit assignment for multi-step reasoning tasks and outperforming existing approaches in fields like mathematics and agent control. Liu et al. [339] propose Direct Advantage Policy Optimization (DAPO), a novel offline RL method that leverages a separately trained critic to evaluate the accuracy of each reasoning step. This technique provides dense feedback for policy optimization, addressing both sparse rewards and training instability. Further, some research focuses on adjusting the focus of RL algorithms to optimize exploration in targeted aspects. Specifically, CPL [570], cDPO [325], Focused-DPO [733] and RFTT [735] enhance exploration in Long CoT by prioritizing critical or error-prone areas through preference optimization, improving accuracy in those regions. Li et al. [300] introduce Learning Impact Measurement (LIM), an automated method for evaluating and prioritizing training samples based on their alignment with model learning trajectories. This approach enables efficient resource use and scalable implementation. For instance, ThinkPO [659] uses short CoT reasoning outputs as rejected answers and longer ones as chosen answers for the same question, applying DPO to encourage prioritization of longer reasoning outputs.
@@
(2) Reward-model-based RL: Earlier, Proximal Policy Optimization (PPO) was first introduced by Schulman et al. [460], which alternates between interacting with the environment to collect data and optimizing a surrogate objective function via stochastic gradient ascent, surpassing DPO [207]. Subsequently, ReMax [311] eliminates the need for additional value models in PPOs. By incorporating variance reduction and REINFORCE [494] techniques, it reduces over four hyperparameters, resulting in lower GPU memory usage and faster training. Building on this, DeepSeekMath [466] proposes Group Relative Policy Optimization (GRPO), replacing traditional value models with improved sampling strategies, thus significantly accelerating learning and achieving performance on par with GPT-4 in mathematics. Hu [181] further refine GRPO with REINFORCE++, simplifying the algorithm and enhancing its training. Additionally, Vassoyan et al. [537] and [784] improve exploration efficiency in smaller models by modifying the KL penalty, thus enhancing performance under distribution shifts. Huang et al. [188] introduce Decoupled Value Policy Optimization (DVPO), a streamlined framework that replaces reward modeling with a pretrained global value model (GVM) and eliminates the interdependence between actor and critic. To address the high-quality demands of reward models, Cui et al. [97] propose PRIME (Process Reinforcement through IMplicit rEwards), which integrates the SFT model as a PRM within a unified reinforcement learning framework, enabling online updates through policy rollouts and outcome labels via implicit process rewards. Finally, Yi et al. [680] introduce SPPD, which employs Process Preference Learning with Dynamic Value Margin for self-training.
@@
DeepSeek-R1-Distill-Llama-70B [155] - - 70.0 - - 57.9
DeepSeek-R1-Distill-Qwen-32B [155] - - 72.6 - - 54.6
@@ 6.2.2 Reward Strategies
The studies explore advancements in training advanced RLLMs using rule-rewarded RL to enhance exploration strategies and reasoning accuracy. These efforts primarily focus on three types of rewards: (1) Correctness Rewarding: Correctness rewards are fundamental for guiding RLLMs toward accurate answers. Specifically, Singh et al. [475] introduce a binary reward system (positive or negative) to facilitate exploration, achieving simple yet effective performance improvements. Similarly, the DeepSeek-R1 [155] employs rule-extracted accuracy as an RL reward, scaling this approach to larger scenarios and training sizes, thereby enhancing both exploration and reasoning tasks [362, 115]. Furthermore, O1-CoderZhang et al. [753], StepCoder [107], and SWE-RL [596] address challenges in code generation by developing a test case generator, which standardizes code testing, ensuring accurate generation. (2) Format Rewarding: Further, format rewards are used to encourage better reasoning paradigms. Guo et al. [155] introduce this concept to effectively guide reasoning and exploration [622]. Xie et al. [622] expanded on this with a three-stage, rule-based RL approach, enabling the Qwen-7B model to learn complex multi-path exploration, which significantly improved both output format and corresponding length consistency. (3) Scaling rewarding: Moreover, scaling rewards are applied to promote longer reasoning chains and broader exploration. Recent studies [64, 411, 243] highlight the need for progressively scaled reasoning lengths to overcome the limitations of current reasoning approaches. As a result, research has focused on scaling exploration [622, 674]. However, excessive scaling can lead to inefficiency and overcomplicated reasoning [96]. Kimi-K1.5 [508] and Arora and Zanette [17] suggest that favoring shorter, more accurate reasoning may also significantly improve efficiency and performance.
@@
It refers to a class of techniques in which RL algorithms are enhanced by leveraging additional reward models, to guide exploration and improve decision-making processes. Earlier in 2021, OpenAI [95] propose a “Gen-Verifier” paradigm to train a correctness-oriented ORM and used ORM-rewarded RL to surpass SFT performance. Recently, with rapid advancements in PRM, several studies [540, 725, 359] have scaled reinforcement learning by enhancing exploration through step-level correctness rewarding. Building on this, Hou et al. [180] introduce entropy rewards and dynamic regularization to further optimize the reasoning process. STeCa [551] identifies suboptimal actions during exploration by comparing step-level rewards and adjusting trajectories to improve deep reasoning. Additionally, the Kimi-K1.5 model [508] extends PRM paradigms into multimodal scenarios, achieving state-of-the-art performance in multi-modal reasoning tasks through a streamlined reinforcement learning framework.
@@ 6.3.1 Human-driven Exploration
Human-driven exploration refers to human-designed constant pipeline exploration for long-term exploration. Several studies highlight the effectiveness of prompt-based [234, 523, 143], tree-structured [780, 668, 67, 441, 388, 33] and even graph-structured [32, 520, 430, 46] search frameworks, demonstrating superior performance and scalability over traditional methods across various datasets. Building on this, CodeTree [284] and Tree-of-Code [396] integrate a tree-based structure with execution and LLM feedback, utilizing multi-agents to optimize multi-stage decisions, thereby improving both strategy planning and solution refinement. Cheng et al. [82] generalize this approach with the Self-Play with Tree-Search Refinement (SPAR) strategy, which generates valid, comparable preference pairs to enhance instruction-following capabilities. Bi et al. [37] and Light et al. [318] extend tree search to a multi-tree paradigm, introducing the Forest-of-Thought framework, which incorporates multiple reasoning trees to improve exploration capabilities to solve complex tasks with greater accuracy.
@@ 6.3.2 Model-driven Exploration
Recent efforts have focused on improving exploration structures during iterations for better logical quality. (1) Beam Search: Earlier, Xie et al. [624] introduced a decoding algorithm that integrates self-evaluation guidance via stochastic beam search, using it as a more reliable automatic criterion to streamline the search in the reasoning space, thereby enhancing prediction quality. Similarly, Zhu et al. [795] propose Deductive Beam Search (DBS), which combines CoT and deductive reasoning with stepwise beam search for RLLMs. (2) A* Search: On another front, Lehnert et al. [269] present Searchformer, which predicts A* algorithm dynamics to improve task performance and reduce search steps [71]. Later, Kang et al. [233] introduce the MindStar (M*) framework, which optimizes reasoning paths through beam search and Levin tree search methods, further enhancing reasoning performance. (3) MCTS Search: Building on the advantages of MCTS, a series of studies, such as Macro-o1 [765], STILL-1 [222], SRA-MCTS [630], utilize MCTS to guide more effective exploration [731, 294, 230, 220, 773, 433, 414]. Xu [634] utilizes energy function for better exploration during Long CoT. Yao et al. [666] further advance this by introducing Collective MCTS (CoMCTS), which leverages collective learning across multiple LLMs to enhance reasoning. Further, MC-NEST [443] integrates Nash Equilibrium strategies to balance exploration and exploitation, improving LLM decision-making in multi-step mathematical tasks. Additionally, CoAT [405] expands the MCTS algorithm with a dynamic correlation memory mechanism, enabling the system to dynamically store new information during inference. Despite MCTS’s benefits, it is often hindered by a large action space and inefficient search strategies, which complicate the generation of Long CoTs. To address this, Lin et al. [322] propose constraining the action space and refining the search strategy to facilitate the emergence of long CoTs. Finally, these methods have been extended to interactive environments, significantly improving success rates in automated exploration tasks [547, 249, 317, 628, 718, 412].

### v3

@@ 5.2.1 Prompt-based Refinement Generation
Research on prompt-based refine generation focuses on enhancing the performance of LLMs through iterative self-refinement mechanisms [458, 864, 82, 374, 820, 604, 654]. A prominent approach involves prompting RLLMs to generate initial outputs, followed by self-feedback that iteratively refines and improves performance across tasks such as dialogue generation and mathematical reasoning [515, 425, 870, 530, 434, 274, 600, 383], which even much reduce the hallucinations [225, 247]. Noteworthy methods, like Self-Backtracking [748], Refiner [469], and BackMath [839], allow LLMs to adjust their reasoning autonomously, reducing unnecessary complexity in decision-making [688]. Further, Havrilla et al. [189] extend the paradigm by integrating overall-level and step-level refinements, improving refinement performance. Yang et al. [751] propose a method to decompose the self-correction capability of LLMs into "confidence" and "critique" capacities, designing probabilistic metrics to evaluate them and exploring the role of reflection mechanisms in model behavior. Additionally, MCTSr [823], LLM2 Yang et al. [736], ReST-MCTS* [822] and ReARTeR [554] emphasize dynamic reflection through iterative error correction and confidence adjustments, allowing models to autonomously refine reasoning strategies [144]. He et al. [191] extend this paradigm to multi-agent scenarios, improving agent system performance [742]. Moreover, Yuksekgonul et al. [800] and Peng et al. [471] further expand the paradigm by enabling automatic prompt optimization driven by LLMs. This approach facilitates more generalized and automated refinement of input prompts across a range of tasks, as opposed to focusing solely on refining output results. However, without oracle feedback, RLLM’s self-refinement process fails, causing instability in both intermediate and final answers, leading to biases in simple factual queries and introducing cognitive biases in complex tasks [836, 721].
@@ 6.1.1 Vertical Scaling
Vertical scaling refers to extending the reasoning output within a single model generation, significantly boosting model performance [308, 837]. Early works by Fu et al. [146] and Jaech et al. [240] show that increasing the length of the reasoning path can greatly improve performance. Tian et al. [587] enhances model reasoning iteratively by using prior answers as prompts for each successive round, thus enabling vertical scaling of the reasoning process. Building on this, later studies [246, 315] further explore enhancing logical depth through tree-based searches within a fixed compute budget, resulting in notable performance gains. Building upon this, Muennighoff et al. [441] introduce a test-time scaling method that improves reasoning by fine-tuning and budget forcing, yielding substantial gains with additional computing at test time. To address the constraints of attention spans, some studies focus on expanding reasoning length in latent spaces. Geiping et al. [159] and Chen et al. [93] enhance test-time reasoning performance by implicitly scaling computation in latent space through recurrent depth.
@@ 6.1.2 Parallel Scaling
Parallel scaling refers to the process of increasing the number of reasoning iterations during model generation and then verfiy these results to get the final output, which significantly enhances model performance [2, 686, 50, 385]. Initially, Wang et al. [652] introduce the concept of self-consistency, demonstrating that multiple sampling processes followed by majority voting for effective exploration.
@@
The primary focus of recent research is optimizing verification, which can be categorized into two types: (1) Overall Verification: Recent works [887, 665] divide the scaling process into two stages: "reasoning" and "self-verification." By replacing majority voting in self-consistency with self-verification, these approaches show significant improvements [859, 70, 907, 291]. In code scenarios, WoT [851], CISC [566] and S* [316] scale the Long CoT in parallel, using output confidence or code execution results for verification, effectively assessing reasoning quality [505, 158, 217, 898]. Further, Nye et al. [449] and Weir et al. [671] train RLLMs to simulate code execution, removing the need for test cases in code-related parallel scaling. Chain-of-Verification [80] introduces meta-verification, sampling multiple verification instances to identify the correct one. Kim et al. [279], Chen et al. [95], and Vacareanu et al. [600] validate this approach empirically by evaluating answer correctness based on reasoning path properties. Moreover, Li et al. [340] tune a specific RLLM to verify and aggregate answers, showing improved performance. This suggests that PRM cannot replace a specially trained RLLM for verification due to training goal biases [856]. Finally, Kang et al. [270] leverage self-uncertainty to select the best results. (2) Step Verification: Building on this, numerous researchers have explored step-level or finer-grained verification [72, 366]. Notably, DIVERSE [343], SSC-CoT [868], and Fine-grained Self-Consistency [80] combine diverse reasoning paths with step-level verification. In addition, Snell et al. [536], Wu et al. [686], Luo et al. [407], Wang et al. [618], Wu et al. [679], and Liu et al. [386] investigate how optimal scaling strategies based on MCTS can enhance smaller language models’ performance. Their findings show that a 1B RLLM can outperform a 405B model on complex tasks through parallel scaling [783]. Despite these advancements in verification, Chen et al. [80] demonstrate that these strategies cannot surpass Best-of-N methods, suggesting that breakthroughs cannot solely rely on optimization-based verification [90].
@@
Another key area of research focuses on generating diverse but less paths or strategies for efficient scaling [691, 614, 69]. For instance, Zeng et al. [812] aggregate the shortest yet most varied reasoning paths for better scalability. Similarly, Du et al. [131] adjust the sampling temperature to increase diversity, leading to improved scaling. Zhang et al. [832] and Liu et al. [375] optimize both candidate solution generation (e.g., prompts, temperature, and top-p) and reward mechanisms (such as self-evaluation and reward types), offering diverse strategies for parallel scaling. Moreover, Qin et al. [490], Luo et al. [410], and Yu et al. [784] enhance RLLM reasoning by scaling sampling across multiple natural and programming languages or varied expressions. Finally, Yang et al. [747] introduces a method where a small set of seed data, with varied response lengths, guides the model to engage in deeper reasoning by selecting the shortest correct responses across various inference efforts.
@@ 6.2.1 RL Strategies
Recent advancements in RL strategies for exploration have led to notable improvements in various tasks, particularly in reasoning tasks [551, 296, 245, 428, 698, 809, 781, 210, 801, 498, 119].
@@
(1) Reward-model-free RL: The first series of work focuses on RL optimization algorithms. Additionally, OREO [621] propose an offline RL method that optimizes the soft Bellman equation, improving credit assignment for multi-step reasoning tasks and outperforming existing approaches in fields like mathematics and agent control. Liu et al. [380] propose Direct Advantage Policy Optimization (DAPO), a novel offline RL method that leverages a separately trained critic to evaluate the accuracy of each reasoning step. This technique provides dense feedback for policy optimization, addressing both sparse rewards and training instability. Further, some research focuses on adjusting the focus of RL algorithms to optimize exploration in targeted aspects. Specifically, CPL [639], cDPO [364], and Focused-DPO [831] enhance exploration in Long CoT by prioritizing critical or error-prone areas through preference optimization, improving accuracy in those regions. Bartoldson et al. [37] further adjusts the replay strategy of the training data, aiming to optimize reasoning performance. Li et al. [339] introduce Learning Impact Measurement (LIM), an automated method for evaluating and prioritizing training samples based on their alignment with model learning trajectories. This approach enables efficient resource use and scalable implementation. For instance, ThinkPO [746] uses short CoT reasoning outputs as rejected answers and longer ones as chosen answers for the same question, applying DPO to encourage prioritization of longer reasoning outputs [896].
@@
(2) Reward-model-based RL: Reward-model-based RL refers to approaches that use a reward model to guide learning and decision-making in the absence of explicit rewards [833]. Earlier, Proximal Policy Optimization (PPO) was first introduced by Schulman et al. [517], which alternates between interacting with the environment to collect data and optimizing a surrogate objective function via stochastic gradient ascent, surpassing DPO [239]. Subsequently, ReMax [350] eliminates the need for additional value models in PPOs. By incorporating variance reduction and REINFORCE [555] techniques, it reduces over four hyperparameters, resulting in lower GPU memory usage and faster training. Building on this, DeepSeekMath [524] proposes Group Relative Policy Optimization (GRPO), replacing traditional value models with improved sampling strategies, thus significantly accelerating learning and achieving performance on par with GPT-4 in mathematics. Hu [207] further refine GRPO with REINFORCE++, simplifying the algorithm and enhancing its training. Additionally, Vassoyan et al. [602] and [888] improve exploration efficiency in smaller models by modifying the KL penalty, thus enhancing performance under distribution shifts. Huang et al. [216] introduce Decoupled Value Policy Optimization (DVPO), a streamlined framework that replaces reward modeling with a pretrained global value model (GVM) and eliminates the interdependence between actor and critic. To address the high-quality demands of reward models, Cui et al. [115] propose PRIME (Process Reinforcement through IMplicit rEwards), which integrates the SFT model as a PRM within a unified reinforcement learning framework, enabling online updates through policy rollouts and outcome labels via implicit process rewards. Finally, Yi et al. [770] introduce SPPD, which employs Process Preference Learning with Dynamic Value Margin for self-training.
@@
SFT Strategies
DeepSeek-R1-Distill-Llama-70B [179] - - 70.0 - - 57.9
DeepSeek-R1-Distill-Qwen-32B [179] - - 72.6 - - 54.6
START [312] QwQ-32B-preview [581] - 66.7 94.4 63.6 47.3
@@ 6.2.2 Reward Strategies
The studies explore advancements in training advanced RLLMs using rule-rewarded RL to enhance exploration strategies and reasoning accuracy. These efforts primarily focus on three types of rewards: (1) Correctness Rewarding: Correctness rewards are fundamental for guiding RLLMs toward accurate answers. Specifically, Singh et al. [534] introduce a binary reward system (positive or negative) to facilitate exploration, achieving simple yet effective performance improvements. Similarly, the DeepSeek-R1 [179] employs rule-extracted accuracy as an RL reward, scaling this approach to larger scenarios and training sizes, thereby enhancing both exploration and reasoning tasks [411, 136]. Furthermore, O1-CoderZhang et al. [854], StepCoder [128], and SWE-RL [670] address challenges in code generation by developing a test case generator, which standardizes code testing, ensuring accurate generation [708, 788]. (2) Format Rewarding: Further, format rewards are used to encourage better reasoning paradigms. Guo et al. [179] introduce this concept to effectively guide reasoning and exploration [701]. Xie et al. [701] expanded on this with a three-stage, rule-based RL approach, enabling the Qwen-7B model to learn complex multi-path exploration, which significantly improved both output format and corresponding length consistency. (3) Scaling rewarding: Moreover, scaling rewards are applied to promote longer reasoning chains and broader exploration. Recent studies [78, 462, 277] highlight the need for progressively scaled reasoning lengths to overcome the limitations of current reasoning approaches. As a result, research has focused on scaling exploration [701, 762]. However, excessive scaling can lead to inefficiency and overcomplicated reasoning [114]. Kimi-K1.5 [572], Yang et al. [747] and Arora and Zanette [19] proposed Long2Short techniques, favoring shorter, more accurate reasoning may also significantly improve efficiency and performance.
@@
It refers to a class of techniques in which RL algorithms are enhanced by leveraging additional reward models, to guide exploration and improve decision-making processes [547]. Earlier in 2021, OpenAI [113] propose a “Gen-Verifier” paradigm to train a correctness-oriented ORM and used ORM-rewarded RL to surpass SFT performance. Recently, with rapid advancements in PRM, several studies [605, 822, 408] have scaled reinforcement learning by enhancing exploration through step-level correctness rewarding [525]. Building on this, Hou et al. [206] introduce entropy rewards and dynamic regularization to further optimize the reasoning process. STeCa [617] identifies suboptimal actions during exploration by comparing step-level rewards and adjusting trajectories to improve deep reasoning. Additionally, the Kimi-K1.5 model [572] extends PRM paradigms into multimodal scenarios, achieving state-of-the-art performance in multi-modal reasoning tasks through a streamlined reinforcement learning framework.
@@ 6.3.1 Human-driven Exploration
Human-driven exploration refers to human-designed constant pipeline exploration for long-term exploration. Several studies highlight the effectiveness of prompt-based [268, 588, 167], tree-structured [884, 755, 81, 496, 438, 43] and even graph-structured [42, 584, 485, 56, 847] search frameworks, demonstrating superior performance and scalability over traditional methods across various datasets. Building on this, CodeTree [323] and Tree-of-Code [446] integrate a tree-based structure with execution and LLM feedback, utilizing multi-agents to optimize multi-stage decisions, thereby improving both strategy planning and solution refinement. Cheng et al. [99] generalize this approach with the Self-Play with Tree-Search Refinement (SPAR) strategy, which generates valid, comparable preference pairs to enhance instruction-following capabilities. Bi et al. [47] and Light et al. [357] extend tree search to a multi-tree paradigm, introducing the Forest-of-Thought framework, which incorporates multiple reasoning trees to improve exploration capabilities to solve complex tasks with greater accuracy. Furthermore, Li et al. [312] explores the integration of Python tools into Long CoT frameworks by both prompting and training, performing test-time scaling more effectively.
@@ 6.3.2 Model-driven Exploration
Recent efforts have focused on improving exploration structures during iterations for better logical quality. (1) Beam Search: Earlier, Xie et al. [703] introduced a decoding algorithm that integrates self-evaluation guidance via stochastic beam search, using it as a more reliable automatic criterion to streamline the search in the reasoning space, thereby enhancing prediction quality. Similarly, Zhu et al. [902] propose Deductive Beam Search (DBS), which combines CoT and deductive reasoning with stepwise beam search for RLLMs. (2) A* Search: On another front, Lehnert et al. [304] present Searchformer, which predicts A* algorithm dynamics to improve task performance and reduce search steps [85]. Later, Kang et al. [267] introduce the MindStar (M*) framework, which optimizes reasoning paths through beam search and Levin tree search methods, further enhancing reasoning performance. (3) MCTS Search: Building on the advantages of MCTS, a series of studies, such as Macro-o1 [867], STILL-1 [255], SRA-MCTS [710], and RFTT [833], utilize MCTS to guide more effective exploration [828, 333, 264, 253, 877, 488, 465]. Xu [714] utilizes energy function for better exploration during Long CoT. Yao et al. [753] further advance this by introducing Collective MCTS (CoMCTS), which leverages collective learning across multiple LLMs to enhance reasoning. Further, MC-NEST [499] integrates Nash Equilibrium strategies to balance exploration and exploitation, improving LLM decision-making in multi-step mathematical tasks. Additionally, CoAT [455] expands the MCTS algorithm with a dynamic correlation memory mechanism, enabling the system to dynamically store new information during inference. Despite MCTS’s benefits, it is often hindered by a large action space and inefficient search strategies, which complicate the generation of Long CoTs. To address this, Lin et al. [361] propose constraining the action space and refining the search strategy to facilitate the emergence of long CoTs. Finally, these methods have been extended to interactive environments, significantly improving success rates in automated exploration tasks [613, 283, 356, 707, 815, 463].
@@
AM-DeepSeek-R1-Distilled [860] Mathematics, Code, General Reward Model + Rule & LLM Validation Lang 1.4M